# Basis Selection: Low-Rank Decomposition of Pretrained Large Language Models for Target Applications

### v2

@@ 4.1 Evaluation Methodology
The performance of low-rank compression algorithms is evaluated on two tasks: mathematical reasoning and code generation. For each task, Llama 2-7B and Llama 2-13B models (Touvron et al., 2023b) are first finetuned on a training dataset and then compressed using a compression algorithm. The compressed models are further finetuned before evaluation.
@@
For the mathematical reasoning task, we utilize two evaluation datasets: GSM8K (Cobbe et al., 2021) and Hendrycks’ MATH (Hendrycks et al., 2021). The GSM8K dataset comprises verbally described mathematical questions, containing 1,319 samples used for evaluation. The Hendrycks’ MATH dataset covers more complex topics such as linear algebra and geometry, consisting of 5,000 question-answer pairs used for evaluation. Due to its complexity, the Hendrycks’ MATH dataset necessitates more sophisticated computations and reasoning, resulting in lower accuracy compared to GSM8K.
For the code generation task, we use two evaluation datasets: MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021a). Both datasets evaluate the models’ ability to generate Python code. MBPP comprises 500 code generation questions, while HumanEval includes 164 questions.
@@
We evaluate our proposed Basel method against state-of-the-art low-rank compression techniques, including SVD and FWSVD. SVD has been widely used in prior work on model compression (Sharma et al., 2024; Acharya et al., 2019; Noach & Goldberg, 2020; Xue et al., 2013; Jaderberg et al., 2014; Denton et al., 2014; Zhang et al., 2015; Povey et al., 2018), and in our experiments, it is applied to compress a model that has been fine-tuned on the target task. FWSVD (Hsu et al., 2022) extends SVD by incorporating gradient-based profiling: it estimates weight importance using gradients computed on the fine-tuning dataset of the target task and leverages this information during decomposition. In addition to low-rank methods, we also compare Basel with leading approaches from other compression paradigms, including QLoRA (Dettmers et al., 2023), a quantization-based method, and FLAP (An et al., 2024), a pruning-based method. QLoRA improves performance by training LoRA adapters on the target task’s fine-tuning dataset. FLAP prunes weights based on their magnitudes and the variance of each input activation dimension, profiled on the target task.
4.2 Evaluating Low-Rank Compression Methods for Mathematical Reasoning
Figures 3 (a) and (b) depict the performance of various low-rank compression algorithms on the Llama 2-7B model for the mathematical reasoning task. We evaluate the models’ accuracy (Pass@1) at different compression ratios (original model size vs. compressed model size). For low compression ratios (below 6), all methods achieve similar accuracy. However, our Basel method significantly outperforms SVD and FWSVD at higher compression ratios. For instance, at a 7x compression ratio, Basel achieves around 46% and 12% accuracy on GSM8K and MATH datasets, respectively, while FWSVD drops below 2% accuracy on both datasets and SVD reaches only 27% and 5% accuracy on GSM8K and MATH, respectively. We also find that Basel, at a compression ratio of 16, achieves better accuracy on both GSM8K and MATH compared to FWSVD and SVD at a compression ratio of 6. This suggests that Basel reduces the model size by up to 2.7 times more than the baseline methods while maintaining similar accuracy. This highlights the effectiveness of Basel for deep compression, especially when aiming for aggressive model size reduction.
@@ 4.2 Evaluating Low-Rank Compression Methods for Mathematical Reasoning
Рис. 3: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms on the mathematical reasoning task.
@@
Рис. 4: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms on the mathematical reasoning task.
@@
Рис. 5: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms on the code generation task.
@@
Рис. 6: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms on the code generation task.
4.3 Evaluating Low-Rank Compression Methods for Code Generation
Similar results extend to code generation tasks (Figures 5 and  6). For both Llama 2-7B and Llama 2-13B models, all methods perform comparably at lower compression ratios (below 4). However, Basel exhibits clear superiority at higher compression ratios. On Llama 2-7B at a 6x compression ratio, Basel achieves 12% and 9% accuracy on HumanEval and MBPP datasets, respectively, significantly outperforming SVD (5% and 2%) and FWSVD (6% and 2%). Similar trends hold for Llama 2-13B. These findings further solidify Basel’s effectiveness for deep compression across diverse tasks and model sizes.
@@ 4.3 Evaluating Low-Rank Compression Methods for Code Generation
Рис. 7: Pass@1 accuracy and model size of Llama 2-7B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the mathematical reasoning task.
@@
Рис. 8: Pass@1 accuracy and model size of Llama 2-13B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the mathematical reasoning task.
4.4 Combining Low-Rank Compression and Quantization for Math Reasoning
@@
To address this, we explore combining quantization with low-rank compression using our Basel method. Figures 7 and 8 illustrate that combining Basel with 8-bit quantization not only achieves a smaller model size but also significantly outperforms 4-bit quantization in accuracy. On compressing Llama 2-7B, Basel + 8-bit achieves 59.6% on GSM8K and 18.1% on MATH with a 3.28 GB model size, compared to 4-bit quantization’s 39.2% and 8.8% with a larger 3.37 GB model. On compressing Llama 2-13B, Basel + 8-bit yields 61.9% on GSM8K and 17.5% on MATH with a 6.13 GB size, while 4-bit quantization achieves only 52.8% and 12.5% with 6.51 GB.
We further compare Basel + 8-bit quantization against two additional approaches: (1) QLoRA with 4-bit quantization and (2) combining 8-bit quantization with FLAP, a state-of-the-art pruning technique. As shown in Figures 7 and 8, Basel + 8-bit consistently outperforms both in terms of accuracy and size. For instance, on compressing Llama 2-7B, it improves accuracy by 5.5% on both GSM8K and MATH compared to QLoRA, while reducing the model size by an additional 0.13 GB.
@@
Рис. 9: Pass@1 accuracy and model size of Llama 2-7B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the code generation task.
@@
Рис. 10: Pass@1 accuracy and model size of Llama 2-13B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the code generation task.
@@ 4.5 Combining Low-Rank Compression and Quantization for Code Generation
We also compare Basel + 8-bit quantization to three alternative methods on the code generation task: (1) standard 4-bit quantization, (2) QLoRA with 4-bit quantization, and (3) FLAP combined with 8-bit quantization. Figures 9 and 10 present the results for compressing Llama 2-7B and Llama 2-13B. On both models, Basel + 8-bit quantization consistently outperforms 4-bit quantization while achieving smaller model sizes. For example, on Llama 2-7B, it improves accuracy by 12.8% on HumanEval and 6.4% on MBPP, with a model size similar to that of the 4-bit version. This further highlights the benefit of combining low-rank compression with quantization.
Basel + 8-bit quantization also significantly outperforms both QLoRA (4-bit) and FLAP + 8-bit quantization on Llama 2-7B and Llama 2-13B. For instance, on compressing Llama 2-7B, it achieves 6.7% higher accuracy on HumanEval and 1.0% higher on MBPP compared to QLoRA, while also reducing the model size by an additional 0.17 GB.
@@ 4.6 Inference
Figure 11 presents the inference throughput and memory consumption of models compressed from Llama-7B on a single A100 GPU, using GSM8K as the evaluation set. The results show that low-rank compression methods, including SVD, FWSVD, and Basel, lead to reduced memory consumption and improved throughput as the model size decreases. Throughput and memory usage are primarily dependent on model size, with no significant differences between the methods at equivalent sizes. However, since our proposed Basel method achieves a greater reduction in model size while maintaining similar accuracy to SVD and FWSVD, it improves throughput by up to 19% and reduces memory consumption by up to 37%.
Рис. 11: Throughput and memory consumption of compressed models.
4.7 Ablation Study
[…]
@@ 4.7 Ablation Study
Рис. 12: Ablation study: Impact of varying the additional dimension in Basel on the compression of Llama 2-7B for the mathematical reasoning task.
@@
Рис. 13: Ablation study: Impact of varying the pruning times in Basel on the compression of Llama 2-7B for the mathematical reasoning task.
@@
Рис. 14: Ablation study: Pass@1 accuracy and size of Llama 2-7B compressed by Basel (proposed) and Basel (free basis) on mathematical reasoning.
Similarly, performing light pruning multiple times allows the model to gradually adapt to parameter reduction, which is especially helpful for extreme compression. Figure 13 illustrates this effect by comparing two pruning strategies—100 times versus 2. The model pruned 100 times consistently achieves higher accuracy at compression ratios greater than 4.
Basel freezes the basis vectors and updates only the singular values (and any additional bases) during compression. To evaluate the benefit of this design choice, we compare it with a variant of Basel where the basis vectors are also updated. Figure 14 presents the performance of both approaches on compressing Llama 2-7B for mathematical reasoning. Freezing the bases results in better accuracy. It also reduces compression time by 33%, as backpropagation involves fewer trainable parameters.
@@
In many real-world LLM deployment scenarios, compute and hardware constraints impose a target model size tailored to a specific application, which may not match any existing pretrained models. For example, Llama-2 offers 7B, 13B, and 70B variants, but no 4B option. To obtain a model of the required size for the target application, one can either (1) train a new model from scratch or (2) compress a nearby larger model. The first approach is often impractical due to high computational costs—training Llama reportedly requires up to one million A100 GPU hours Touvron et al. (2023a)—and limited access to pretraining data. In contrast, model compression is far more efficient; low-rank methods such as Basel typically require less than 0.01% of the pretraining compute.
Таблица 2: Pass@1 accuracy and model size of Llama 3.2-3B compression via Basel for mathematical reasoning.

### v4

@@ 4.1 Evaluation Methodology
We evaluate the performance of low-rank compression algorithms on three tasks: mathematical reasoning, code generation, and language modeling. For the mathematical reasoning task, we utilize two evaluation datasets: GSM8K (gsm8k) and Hendrycks’ MATH (hendrycks). The GSM8K dataset comprises verbally described mathematical questions, containing 1,319 samples used for evaluation. The Hendrycks’ MATH dataset covers more complex topics such as linear algebra and geometry, consisting of 5,000 question-answer pairs used for evaluation. Due to its complexity, the Hendrycks’ MATH dataset necessitates more sophisticated computations and reasoning, resulting in lower accuracy compared to GSM8K. For the code generation task, we use two evaluation datasets: MBPP (mbpp) and HumanEval (humaneval). Both datasets evaluate the models’ ability to generate Python code. MBPP comprises 500 code generation questions, while HumanEval includes 164 questions. For the language modeling task, we evaluate on WikiText-2 (wikitext2), a dataset consisting of Wikipedia articles.
@@
We evaluate Basel against low-rank compression methods—SVD, FWSVD, and SVD-LLM—as well as approaches from other compression paradigms. SVD, a widely used technique in prior work on model compression (LASER; SVD_recent2; SVD_recent3; SVD0; SVD1; SVD2; SVD3; SVD4), is applied in our experiments to compress models that have already been fine-tuned on the target task. FWSVD (FWSVD) extends SVD by incorporating gradient-based profiling: it estimates weight importance using gradients computed on the fine-tuning dataset and leverages this information during decomposition. SVD-LLM (svdllm) reconstructs weight matrices in low-rank form by minimizing the discrepancy between activations before and after decomposition. Beyond low-rank approaches, we also compare Basel with quantization and pruning methods. QLoRA (QLoRA) is a quantization-based technique that improves performance by training LoRA adapters on the fine-tuning dataset. FLAP (FLAP) and Wanda (wanda) are pruning-based methods that prune weights based on weight magnitudes, together with either the magnitude (Wanda) or the variance (FLAP) of input activations profiled on the target task.
Our implementation of Basel is configured with the following key hyperparameters: KeepRatio varies from 70% to 5%, PruningTimes = 100, KeepingEpoch = 1, PruningEpoch = 2 (for math reasoning and code generation) and 1 (for language modeling), PostFineTuningEpoch = 3, and $\tilde{r}=32$ (see Algorithm 1 for further details).
@@
Рис. 3: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms and pruning algorithms on the mathematical reasoning task. Exact values are listed in Tables 4 and 5 in the appendix.
@@
Рис. 4: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms and pruning algorithms on the math reasoning task. Exact values are listed in Table 6 and 7 in the appendix.
4.2 Evaluating Low-Rank Compression Methods and Pruning Methods for Mathematical Reasoning
Figures 3 (a) and (b) depict the performance of various low-rank compression algorithms on the Llama 2-7B model (llama2) for the mathematical reasoning task. We evaluate the models’ accuracy (Pass@1) at different compression ratios (original model size vs. compressed model size). For low compression ratios (below 6), all low-rank compression methods achieve similar accuracy. However, our Basel method significantly outperforms SVD and FWSVD at higher compression ratios. For instance, at a 7x compression ratio, Basel achieves around 46% and 12% accuracy on GSM8K and MATH datasets, respectively, while FWSVD drops below 2% accuracy on both datasets and SVD reaches only 27% and 5% accuracy on GSM8K and MATH, respectively. We also find that Basel, at a compression ratio of 16, achieves better accuracy on both GSM8K and MATH compared to FWSVD and SVD at a compression ratio of 6. This suggests that Basel reduces the model size by up to 2.7 times more than SVD and FWSVD while maintaining similar accuracy. This highlights the effectiveness of Basel for deep compression, especially when aiming for aggressive model size reduction.
Figures 3 and 4 also compare Basel with the pruning algorithms FLAP and Wanda on GSM8K and MATH. The results show that Basel achieves substantially greater model size reduction while preserving comparable or superior performance.
@@ 4.2 Evaluating Low-Rank Compression Methods and Pruning Methods for Mathematical Reasoning
Рис. 5: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms and pruning algorithms on the code generation task. Exact values are listed in Tables 8 and 9 in the appendix.
@@
Рис. 6: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms and pruning algorithms on the code generation task. Exact values are listed in Tables 10 and 11 in the appendix.
4.3 Evaluating Low-Rank Compression Methods and Pruning Methods for Code Generation
Similar results extend to code generation tasks (Figures 5 and 6). For both Llama 2-7B and Llama 2-13B models, all low-rank compression methods perform comparably at lower compression ratios (below 4). However, Basel exhibits clear superiority at higher compression ratios. On Llama 2-7B at a 6x compression ratio, Basel achieves 12% and 9% accuracy on HumanEval and MBPP datasets, respectively, significantly outperforming SVD (5% and 2%) and FWSVD (6% and 2%). Similar trends hold for Llama 2-13B. Moreover, Basel consistently outperforms pruning baselines FLAP and Wanda, particularly under aggressive compression. These findings further solidify Basel’s effectiveness for deep compression across diverse tasks and model sizes.
4.4 Evaluating Low-Rank Compression Methods for Language Modeling
For the language modeling task, we compress Llama-7B using Basel and baseline methods SVD, FWSVD, and SVD-LLM, and evaluate the resulting models in terms of perplexity. Perplexity serves as a measure of language modeling quality, with lower values indicating better performance. As shown in Figure 7, Basel consistently yields lower perplexity than SVD, FWSVD, and SVD-LLM at the same compression ratio. Conversely, for a given perplexity, Basel produces a substantially smaller model. For instance, Basel achieves a perplexity of 10.45 at a $10\times$ compression ratio, whereas SVD-LLM reaches a perplexity of 15.00 at only $2.5\times$. This demonstrates that Basel attains up to four times greater compression than SVD-LLM (and even more than SVD and FWSVD) while delivering superior performance.
Рис. 7: Perplexity and model size of Llama-7B compressed with various low-rank algorithms on WikiText-2. Lower perplexity indicates better language modeling performance. Results for SVD, FWSVD, and SVD-LLM are taken from (svdllm). Exact values are listed in Table 12 in the appendix.
@@ 4.4 Evaluating Low-Rank Compression Methods for Language Modeling
Рис. 8: Pass@1 accuracy and model size of Llama 2-7B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the mathematical reasoning task. Exact values are provided in Table 13 in the appendix.
@@
Рис. 9: Pass@1 accuracy and model size of Llama 2-13B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the mathematical reasoning task. Exact values are provided in Table 14 in the appendix.
4.5 Combining Low-Rank Compression and Quantization for Mathematical Reasoning
@@
To address this, we explore combining quantization with low-rank compression using our Basel method. Figures 8 and 9 illustrate that combining Basel with 8-bit quantization not only achieves a smaller model size but also significantly outperforms 4-bit quantization in accuracy. On compressing Llama 2-7B, Basel + 8-bit achieves 59.6% on GSM8K and 18.1% on MATH with a 3.28 GB model size, compared to 4-bit quantization’s 39.2% and 8.8% with a larger 3.37 GB model. On compressing Llama 2-13B, Basel + 8-bit yields 61.9% on GSM8K and 17.5% on MATH with a 6.13 GB size, while 4-bit quantization achieves only 52.8% and 12.5% with 6.51 GB.
We further compare Basel + 8-bit quantization against three additional approaches: (1) QLoRA with 4-bit quantization, (2) combining 8-bit quantization with FLAP, and (3) combining 8-bit quantization with Wanda. As shown in Figures 8 and 9, Basel + 8-bit consistently outperforms both in terms of accuracy and size. For instance, on compressing Llama 2-7B, it improves accuracy by 5.5% on both GSM8K and MATH compared to QLoRA, while reducing the model size by an additional 0.13 GB.
@@
Рис. 10: Pass@1 accuracy and model size of Llama 2-7B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the code generation task. Exact values are provided in Table 15 in the appendix.
@@
Рис. 11: Pass@1 accuracy and model size of Llama 2-13B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the code generation task. Exact values are provided in Table 16 in the appendix.
Рис. 12: Throughput and memory consumption of compressed models.
@@ 4.6 Combining Low-Rank Compression and Quantization for Code Generation
We also compare Basel + 8-bit quantization to three alternative methods on the code generation task: (1) standard 4-bit quantization, (2) QLoRA with 4-bit quantization, (3) FLAP combined with 8-bit quantization, and (4) Wanda combined with 8-bit quantization. Figures 10 and 11 present the results for compressing Llama 2-7B and Llama 2-13B. On both models, Basel + 8-bit quantization consistently outperforms 4-bit quantization while achieving smaller model sizes. For example, on Llama 2-7B, it improves accuracy by 12.8% on HumanEval and 6.4% on MBPP, with a model size similar to that of the 4-bit version. This further highlights the benefit of combining low-rank compression with quantization.
Basel + 8-bit quantization also significantly outperforms both QLoRA (4-bit), FLAP + 8-bit quantization, and Wanda + 8-bit quantization on Llama 2-7B and Llama 2-13B. For instance, on compressing Llama 2-7B, it achieves 6.7% higher accuracy on HumanEval and 1.0% higher on MBPP compared to QLoRA, while also reducing the model size by an additional 0.17 GB.
@@ 4.7 Inference
Figure 12 presents the inference throughput and memory consumption of models compressed from Llama 2-7B on a single A100 GPU, using GSM8K as the evaluation set. The results show that low-rank compression methods, including SVD, FWSVD, and Basel, lead to reduced memory consumption and improved throughput as the model size decreases. Throughput and memory usage are primarily dependent on model size, with no significant differences between the methods at equivalent sizes. However, since our proposed Basel method achieves a greater reduction in model size while maintaining similar accuracy to SVD and FWSVD, it improves throughput by up to 19% and reduces memory consumption by up to 37%.
@@
Рис. 13: Ablation study: Impact of varying the additional dimension in Basel on the compression of Llama 2-7B for the mathematical reasoning task. Exact values are listed in Table 19 in the appendix.
@@
Рис. 14: Ablation study: Impact of varying the pruning times in Basel on the compression of Llama 2-7B for the mathematical reasoning task. Exact values are listed in Table 20 in the appendix.
@@
Рис. 15: Ablation study: Pass@1 accuracy and size of Llama 2-7B compressed by Basel (proposed) and Basel (free basis) on mathematical reasoning. Exact values are listed in Table 21 in the appendix.
(a) GSM8K
(b) MATH
Рис. 16: Ablation study: Pass@1 accuracy and model size of Llama-2-7B compressed with Basel, with and without $L1$ regularization, on the mathematical reasoning task. Exact values are reported in Table 22 in the appendix.
4.8 Ablation Study
We conduct a series of ablation studies to better understand the design of Basel.
Impact of hyperparameters. We first evaluate the effects of two key hyperparameters: the additional dimension ($\tilde{r}$ in Equation (3)) and the number of pruning iterations. The additional dimension is introduced to help recover information lost during pruning, particularly under high compression. As shown in Figure 13, setting $\tilde{r}=32$ improves accuracy when compressing Llama 2-7B for mathematical reasoning, especially beyond a $7\times$ compression ratio. However, this accounts for only a small fraction of Basel’s overall performance gain relative to baselines (Figures 3 and 13). Similarly, applying pruning gradually over many iterations allows the model to better adapt to parameter reduction, which is especially beneficial under extreme compression. Figure 14 illustrates this effect: pruning 100 times consistently yields higher accuracy than pruning only twice when the compression ratio exceeds 4.
Freezing basis vectors. Basel freezes the basis vectors and updates only the singular values (and any additional bases) during compression. To evaluate this design choice, we compare Basel with a variant that also updates the basis vectors. As shown in Figure 15, freezing the bases results in better accuracy on Llama 2-7B for mathematical reasoning. This strategy also reduces compression time by 33%, since backpropagation involves fewer trainable parameters.
Inclusion of $L1$ regularization. Finally, we examine the effect of adding an $L1$ penalty on the learnable singular values $\tilde{s}_{i}$. The goal is to encourage sparsity in $\tilde{s}_{i}$, thereby enabling additional compression. We experiment with two regularization weights, $\lambda=0.02$ and $\lambda=0.1$. Figure 16 reports the results. With $\lambda=0.02$, performance is similar to standard Basel; however, a larger weight ($\lambda=0.1$) degrades performance under deep compression.
@@ 4.8 Ablation Study
In many real-world LLM deployment scenarios, compute and hardware constraints impose a target model size tailored to a specific application, which may not match any existing pretrained models. For example, Llama-2 offers 7B, 13B, and 70B variants, but no 4B option. To obtain a model of the required size for the target application, one can either (1) train a new model from scratch or (2) compress a nearby larger model. The first approach is often impractical due to high computational costs—training Llama reportedly requires up to one million A100 GPU hours llama—and limited access to pretraining data. In contrast, model compression is far more efficient; low-rank methods such as Basel typically require less than 0.01% of the pretraining compute.
Таблица 3: Pass@1 accuracy and model size of Llama 3.2-3B compression via Basel for mathematical reasoning.